OpenStudy (hkmiu):

The locations, given in polar coordinates, for two planes approaching an airport are (4 mi, 12º) and (3 mi, 73º). Find the distance between the two planes. a. 1.00 mi c. 5.00 mi b. 61.01 mi d. 3.66 mi

OpenStudy (anonymous):

the included angle is 73 - 12 = 61 degrees. Now just use law of cosine

OpenStudy (hkmiu):

cosine of 61?

OpenStudy (anonymous):

given a,b and an included angle θ, the length of the missing side, called c, is: c^2 = a^2 + b^2 - 2ab cos(θ)

OpenStudy (anonymous):

you know two sides, and the angle. It's just the matter of plug and chug

OpenStudy (hkmiu):

either im sleepy or I just don't get it at all D:

OpenStudy (anonymous):

you're sleepy :D

OpenStudy (hkmiu):

its about 3.66 mi right